OpenStudy (anonymous):

check all that apply, if tan e= 15/8, then: A: csc e=17/15 B: cos e= 15/17 C: cot e= 8/15 D:sec e=17/8

OpenStudy (anonymous):

definitely C. I haven't done the others though.

OpenStudy (anonymous):

how do you solve this though? i need to learn for a test

OpenStudy (anonymous):

well cot is the opposite of tan\[\tan=\frac{ opp }{ adj } ~~~~\cot=\frac{ adj }{ opp }\]

OpenStudy (anonymous):

what about the rest

OpenStudy (anonymous):

\[\sec=\frac{ 1 }{ \cos }~~so~~\sec=\frac{ hyp }{ adj }\]Do you see how it flips the cos?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

ok. So cosine is \[\frac{ adjacent }{ hypotenuse }\] Secant is the inverse of cosine. So you flip them\[\frac{ adjacent }{ hypotenuse }~~~goes \to~~~\frac{ hypotenuse }{ adjacent }\]Does that make sense? (sorry I'm at work, so I'm not able to explain as fast as I usually do.

OpenStudy (anonymous):

okay so far that makes sense

OpenStudy (anonymous):

what is the hypotenuse for this one though? like how do you find it out

OpenStudy (anonymous):

Ok good deal. Do you know how to use the Pythagorean theorem? That's how you'd find the hypotenuse

OpenStudy (anonymous):

yeah i do so would it be like a^2+ 15^2=8^2 if you do this then you get 12.68 but that doesnt seem right

OpenStudy (anonymous):

Your 8 is in the wrong place. \[\tan=\frac{ opp }{ adj }\]You don't use hypotenuse for tan. \[8^{2}+15^{2}=c\]is what you'd use

OpenStudy (anonymous):

\[c ^{2}\]sorry

OpenStudy (anonymous):

oh wow im stupid! haha thank you

OpenStudy (anonymous):

Haha, no I used to do it all the time. Eventually, you don't even think about which one is which.

OpenStudy (anonymous):

Were you able to figure out the hypotenuse?